The production budget is typically prepared before the direct materials budget. True or False True FalsePetrini Corporation makes one product and it provided the following information to help prepare the master budget for the nexMultiple Choice $20.80 $87.80 $100.60 O $79.80Hesterman Corporation makes one product and has provided the following information to help prepare the master budget for theThe ending finished goods inventory should equal 40% of the following months sales. The ending raw materials inventory shoul
Answer #1

1) Yes, Production budget should be prepared before direct material budget

So above statement is true

2) Unit product cost

Direct material (5*4) 20
Direct labor (23*2.6) 59.8
Overhead (8*2.6) 20.8
Unit product cost 100.60

So answer is c) $100.60

3) Production Unit = 9400+(14000*40%)-(9400*40%) = 11240 Units

Raw material purchase = (11240*3)+(39720*20%)-(11240*3*20%) = 34920

Raw material cost = 34920*3 = 104760

So answer is c) $104760